KSU Corp. is considering purchasing one of two new diagnostic machines. Either machine would make it possible for the company to bid on jobs that it currently isn't equipped to do. Estimates regarding each machine are provided below.

                       Machine A,Machine B

Original cost $106,000 $ 175,000

Estimated life  8 years 8 years

Salvage value$3,000 -0-

Estimated annual cash inflows $30,000 $45,000

Estimated annual cash outflows $10,000 $15,000

(A) Calculate the following for each machine. (i). Net present value assuming a 9% discount rate. (ii). Profitability index(B) Which machine should be purchased? Why?


1
Expert's answer
2021-07-01T05:03:49-0400

A)

i) "NPV_A=-106000+(30000-10000)*\frac{\frac{1}{1.09}*(1-\frac{1}{1.09}^8)}{1-\frac{1}{1.09}}+\frac{3000}{1.09^8}=6201.98"

"NPV_B=-175000+(45000-15000)*\frac{\frac{1}{1.09}*(1-\frac{1}{1.09}^8)}{1-\frac{1}{1.09}}=-8955.43"

ii) "PI_A=1+\frac{6201.98}{106000}=1.059"

"PI_B=1+\frac{-8955.43}{175000}=0.949"

B) I will choose machine A because its NPV is positive and the profit index is greater than 1.